### v1

@@ 5 Discussion and Conclusion
In this work, we show that in the framework of GR, observations of the shift in the detector frame chirp mass of the GW sources in two different frequency bands and inference of the luminosity distance provide a new way to measure the expansion rate of the Universe without requiring the use of any EM observations nor leveraging on any astrophysical assumption of the GW source. Due to the advantage of the GW signal not being susceptible to contamination from foreground contamination such as dust, the measure can be possible up to high redshift, to the epoch of the first ever binary black hole in the Universe.
@@
This new frontier of observational cosmology if realized from the GW detectors, can provide the best possible way to measure the high redshift Universe. However, such measurements will require control of several sources of errors both on the instrument side as well as on the waveform modeling side. This kind of measurement will require waveform accuracy in the inspiral and merger phase up to $10^{-11}$ to successfully infer cosmology from GW signals. The degeneracy between inclination angle and luminosity distance [18], and impact from weak lensing [27] are dominant sources of uncertainties on the luminosity distance side. On the chirp mass inference side, the chirp mass inference can be impacted by environmental effects. As we are only focusing on stellar origin BBHs, these black holes are unlikely to have accretion discs or not be embedded only near a galactic center. As a result, the dephasing of the GW emission due to any local astrophysical effects is negligible.
@@
As we are at an exciting time of designing the next generation of GW detectors, appropriate planning for observational strategies and advanced detector sensitivities can make a paradigm shift in observational cosmology using this technique. This technique brings a way to measure the nature of dark energy which is not limited by astrophysical modeling uncertainties and can be achieved by precise measurement from GW detectors operational from mili-Hertz to hecto-Hertz range. Beyond dark energy, this technique makes it also possible to independently map the expansion rate at high redshift, which can discover new components in the Universe that may be dominant only at high redshift. In summary, GW detector development which can make robust and precise measurements of the chirp mass, can open a new window to fundamental physics. This technique makes GW a self-sufficient probe to explore cosmology without depending on any EM observations.
[…]
@@
Acknowledgements This work is part of the ⟨data|theory⟩ Universe-Lab, funded by TIFR and the Department of Atomic Energy, Government of India. The author acknowledges the use of computer cluster of the ⟨data|theory⟩ Universe-Lab and using Astropy [30, 31], Giant-Triangle-Confusogram [32], IPython [33], Matplotlib [34], NumPy [35], and SciPy [36] in this analysis.

### v2

@@ 6 Discussion and Conclusion
In this work, we show that in the framework of GR, observations of the shift in the detector frame chirp mass of the GW sources in two different frequency bands and inference of the luminosity distance provide a new way to measure the expansion rate of the Universe without requiring the use of any EM observations nor leveraging on any astrophysical assumption of the GW source. Due to the advantage of the GW signal not being susceptible to foreground contamination such as dust, the measurement is possible up to high redshift, to the epoch of the first ever binary black hole in the Universe.
@@
This new frontier of observational cosmology, if realized from the GW detectors, can provide the best possible way to measure the high redshift Universe. However, such measurements will require control of several sources of errors, both on the instrument side as well as on the waveform modeling side. This kind of measurement will require waveform accuracy in the inspiral and merger phase up to $10^{-11}$ to successfully infer cosmology from GW signals. The degeneracy between inclination angle and luminosity distance (Flanagan & Hughes, 1998), and impact from weak lensing (Hirata et al., 2010) are dominant sources of uncertainties on the luminosity distance side. On the chirp mass inference side, the chirp mass inference can be impacted by environmental effects. As we are only focusing on stellar origin BBHs, these black holes are unlikely to have accretion discs or be embedded only near a galactic center. As a result, the dephasing of the GW emission due to any local astrophysical effects is negligible.
@@
As we are at an exciting time of designing the next generation of GW detectors, appropriate planning for observational strategies and advanced detector sensitivities can make a paradigm shift in observational cosmology using this technique. This technique brings a way to measure the nature of dark energy, which is not limited by astrophysical modeling uncertainties and can be achieved by precise measurement from GW detectors operational from mili-Hertz to hecto-Hertz range. Beyond dark energy, this technique also makes it possible to independently map the expansion rate at high redshift, which can discover new components in the Universe that may be dominant only at high redshift. In summary, GW detector development, which can make robust and precise measurements of the chirp mass, can open a new window to fundamental physics. This technique makes GW a self-sufficient probe to explore cosmology without depending on any EM observations.
Data Availability
The data underlying this article will be shared at the request to the corresponding author.
Acknowledgements
@@
This work is part of the ⟨data|theory⟩ Universe-Lab, funded by TIFR and the Department of Atomic Energy, Government of India. The author acknowledges the use of computer cluster of the ⟨data|theory⟩ Universe-Lab and using Astropy (Astropy Collaboration et al., 2013, 2018), Giant-Triangle-Confusogram (Bocquet & Carter, 2016), IPython (Pérez & Granger, 2007), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), and SciPy (Jones et al., 01) in this analysis.